OpenStudy (anonymous):

x(x+12)=-28 I got (x+6)=-28. Is this correct? Please and Thank you.

Directrix (directrix):

@kitsune0724 Are you supposed to solve the given equation for x?

OpenStudy (anonymous):

nope first multiply and then simplify and then

OpenStudy (anonymous):

@Directrix No you are suppose to find what would give the middle of the problem.

Directrix (directrix):

@kitsune0724 I don't know what you mean by "middle of the problem." x(x+12)=-28 x^2 + 12x = -28 x^2 + 12x + 28 = 0 After that, I don't know what to say because I don't know the "middle."

OpenStudy (anonymous):

@Directrix would it be like this (x-12)^2=-28?

Directrix (directrix):

@kitsune0724 I don't know because I don't understand what we are supposed to be doing. We need the directions / instructions for the problem.

OpenStudy (anonymous):

@Directrix Which of the following equations is equivalent to the equation in exercise 30? (In other words, which of these is an equation you should get in the MIDDLE of the completing the square process?) this is the problem.

OpenStudy (anonymous):

Oh...You complete the square...

OpenStudy (anonymous):

He forgot the ^2

Directrix (directrix):

x(x+12)=-28 x^2 + 12x = -28 x^2 + 12x + 36 = -28 + 36 (x + 6)* (x+6) = 8 -----> Do you think we've made it to the middle yet. @kitsune0724

OpenStudy (anonymous):

@Directrix yes. that's what I got so far.

Directrix (directrix):

(x + 6)* (x+6) = 8 (x + 6) = sqrt(8) OR (x + 6) = - sqrt (8) @kitsune0724

OpenStudy (anonymous):

@Directrix (x + 6) = - sqrt (8).

Directrix (directrix):

@kitsune0724 What about this--> (x + 6) = sqrt(8) It is one of the two roots.

OpenStudy (anonymous):

@Directrix would it be x=6+2sqrt2, 6-2sqrt2?

Directrix (directrix):

Not quite. I think you have a sign error. (x + 6) = sqrt(8) OR (x + 6) = - sqrt (8) @kitsune0724 x = -6 + sqrt(8) OR x = -6 - sqrt (8) x = -6 + 2sqrt2 or x = -6 -2sqrt2

OpenStudy (anonymous):

@Directrix x=-6+2sqrt2, -6-2sqrt2? I'm getting a bit confused here. Can you please draw it out for so I can what I'm doing wrong? Thank you.

Directrix (directrix):

We're just subtracting 6 from both sides to solve for x. For example, (x + 6) = - sqrt (8). Subtract 6 from both sides. x+6 - 6 = -sqrt(8) - 6 x + 0 = -6 - sqrt(8) x = -6 - 2sqrt2 Same thinking for the second root. @kitsune0724